Chapter: Arithmetic progression
Grade:10
Name: Devi
School: Gems our own indian school
1+4+7+10+............+x=287 , find the value of x

Grade:10

1 Answers

Vikas TU
4 years ago
Dear student
Here first term = 1 
Common difference = 3 
x = a+(n-1)d 
x = 1 + 3n -3 = 3n – 2 
Sum = 287 
n/2 (a+x) = 287 
n/2 (1+ 3n -2) = 287 
3n^2 / 2  – n/2 = 287 
3n^2 – n – 574 =0 
n = 14 , n = -41/3 
so n = 14 
x = 3n -2 = 42 -2 = 40 
Hope this helps 
Good Luck